consider the probability that no more than 19 out of 158 flights will be delayed. assume the probability that a given flight will be delayed is 13%. approximate the probability using normal distibution.
Answer #1

Answer)

Given P = 0.13 (13%)

N = 158

Mean is given by n*p = 20.54

Standard deviation (s.d) = √n*p*(1-p) = 4.22726862169

We need to find P(x<=19)

By continuity correction we will find

P(x<19.5)

Z = (x-mean)/s.d

Z = (19.5-20.54)/4.22726862169

Z = -0.25

From z table, P(Z<-0.25) = 0.4013